# Quantum teleportation implies symmetry-protected topological order

### v2

@@ 6.2 Hypergraph states
We present a $k=1$ standard teleportation protocol involving a hypergraph Rossi et al. (2013); Gachechiladze et al. (2019); Miller and Miyake (2016) resource state $\lvert\Psi_{t}\rangle$ (36). Because $\operatorname*{stab}\left(*\right){\lvert\Psi_{t}\rangle}\nsubseteq\mathbf{P}_{*}{N}$ for the hypergraph state $\lvert\Psi_{t}\rangle$, this example highlights how our results extend beyond “stabilizer states” Gottesman (1997). The hypergraph state $\lvert\Psi_{t}\rangle$ is defined on the vertices $v\in V$ of a graph $G=(V,E)$ comprising connected tetrahedra, as depicted in Fig. 6. The initial state is the same as for the cluster state (20)
$\lvert\Psi_{0}\rangle=\lvert\psi\rangle_{i}\otimes\lvert+++\cdots\rangle\otimes\lvert\boldsymbol{0}\rangle_{\rm ss}\,,$ (68)
@@
$\mathcal{U}=\prod_{\Delta\in\mathcal{F}_{u}}\mathrm{CCZ}_{\Delta}\,,$ (69)
where $\Delta$ runs over all unshaded faces (the set $\mathcal{F}_{u}$) of the graph in Fig. 6, and the controlled-controlled-$Z$ gate $\mathrm{CCZ}_{\Delta}$ applies CZ to two of the qubits on the face $\Delta$ if the third is in the $Z$-basis state $\lvert 1\rangle$. In the computational ($Z$) basis, $\mathrm{CCZ}_{\Delta}=\mathrm{diag}(1,1,1,1,1,1,1,-1)$ sends $\lvert 111\rangle_{\Delta}\to-\lvert 111\rangle_{\Delta}$, and acts trivially otherwise.
We note that a length-three version of this hypergraph state appears in Ref. 46, but its connection to SPT order was not discussed. The hypergraph state in Fig. 6 can be viewed as a cluster state Briegel and Raussendorf (2001) whose odd sites correspond to the vertices connecting the tetrahedra (noted by dots in Fig. 6), and whose even sites are encoded in the three qubits around the shaded faces in Fig. 6.
The CCZ gate corresponds to the third order of the Clifford hierarchy Gottesman and Chuang (1999), where the first order is the Pauli group $\mathbf{P}_{*}{N}$, the second order is the Clifford group $\mathbf{C}_{*}{N}$, which maps $\mathbf{P}_{*}{N}$ to itself, while gates like CCZ map $\mathbf{C}_{*}{N}$ to itself. In particular, for three vertices $i,j,k$, we have
$\mathrm{CCZ}_{ijk}\,X_{i}\,\mathrm{CCZ}_{ijk}=X_{i}\,\mathrm{CZ}_{jk}\,,~{}$ (70)
@@
The hypergraph teleportation protocol $\mathcal{W}$ is straightforward. After applying the entangling unitary $\mathcal{U}$ (69) to the initial state $\lvert\Psi_{0}\rangle$ (68), one next measures all qubits (except the rightmost site $f$) in the $X$ basis in the channel $\mathcal{M}$. Thus, the measurement regions in Fig. 6 correspond to the tetrahedra formed by the shaded triangular faces and the vertex to the left. In the channel $\mathcal{R}$, the parity of the outcomes of all $X_{j}$ measurements for qubits $j\in\mathcal{F}_{s}$ in the shaded triangles in Fig. 6 determine a final $\chi_{z}=X_{f}$ rotation, while the parity of the $X_{j}$ outcomes for all other qubits $j\in V\setminus\mathcal{F}_{s}$ determine a final $\chi_{x}=Z_{f}$ rotation.
In the Heisenberg picture, the two logical operators $X_{{\rm L}}$ and $Z_{{\rm L}}$ evolve under the dilated channels $\mathcal{R}$ and $\mathcal{M}$ via
$X^{\prime}_{\rm L}$ $=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{M}^{\dagger}\,\mathcal{R}^{\dagger}\,X_{f}\,\mathcal{R}\,\mathcal{M}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=\prod_{j\in V\setminus\mathcal{F}_{s}}X_{j}$ (71a)
$Z^{\prime}_{\rm L}$ $=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{M}^{\dagger}\,\mathcal{R}^{\dagger}\,Z_{f}\,\mathcal{R}\,\mathcal{M}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=Z_{f}\prod_{\Delta\in\mathcal{F}_{s}}X_{\Delta}\,,$ (71b)
acting on the hypergraph resource state $\lvert\Psi_{t}\rangle$, where $\mathcal{F}_{s}$ denotes the shaded faces in Fig. 6, the set $V\setminus\mathcal{F}_{s}$ corresponds to the filled dots in Fig. 6, and $X_{\Delta}=\prod_{j\in\Delta}\,X_{j}$.
@@
$X_{f}(T)$ $=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{W}^{\dagger}\,X_{f}\,\mathcal{W}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=\prod_{j\in V\setminus\mathcal{F}_{s}}X_{j}$ (72a)
$Z_{f}(T)$ $=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{W}^{\dagger}\,Z_{f}\,\mathcal{W}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=Z_{i}\prod_{\Delta\in\mathcal{F}_{s}}X_{\Delta}\,,$ (72b)
where $\mathcal{U}$ acts trivially on all $X$ operators. Essentially, $\mathcal{U}$ (69) acts on $X^{\prime}_{\rm L}$ (71a) by attaching CZ operators to $X_{j}$ along all edges of each shaded face whose vertices neighbor qubit $j$ (70). However, each shaded face borders two such $X_{j}$ operators, and so each CZ operator is attached twice to $X^{\prime}_{\rm L}$ (71a), but $\mathrm{CZ}^{2}=\mathds{1}$. The action of $\mathcal{U}$ (69) on $Z^{\prime}_{\rm L}$ (71b) again follows from Eq. 70, and again, we find that CZ gates are always attached twice, so that the only change from Eq. 71b to Eq. 72b is that $Z_{f}\to Z_{i}$. Note that Eqs. 6.2 satisfy Prop. 1 for the initial state $\lvert\Psi_{0}\rangle$ (68).
Finally, we identify the string order parameters establishing that the hypergraph resource state $\lvert\Psi_{t}\rangle$ has nontrivial $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT order. The string order parameters (53) correspond to the stabilizers $\mathsf{S}_{\nu,t}$ given by
$\mathsf{S}_{\nu,t}$ $=\mathcal{U}\,\mathsf{S}_{\nu}\,\mathcal{U}^{\dagger}=\mathcal{U}\,\sigma^{\nu}_{i}\,\sigma^{\nu}_{f}(T)\,\mathcal{U}^{\dagger}$
@@
which is the product of the initial-state logical operator $\sigma^{\nu}_{i}$ evolved “forward” under $\mathcal{U}$ (69) and the resource-state logical operators (6.2). We note that $Z_{i}$ commutes with $\mathcal{U}$(69), and is thus unchanged, while $X_{i}$ evolves according to Eq. 70. The resulting order parameters are
$\mathsf{S}_{x,t}$ $=\mathrm{CZ}_{\Delta_{i}}\,\Big{(}\prod\limits_{j\in\mathcal{V}}\,X_{j}\,\Big{)}\,X_{f}$ (74a)
$\mathsf{S}_{z,t}$ $=Z_{i}\,X_{\Delta_{i}}\,\Big{(}\prod\limits_{\begin{subarray}{c}\Delta\in\mathcal{F}_{s}\\ \delta\neq\Delta_{i}\end{subarray}}X_{\Delta}\,\Big{)}\,Z_{f}\,,~{}$ (74d)
where $\mathrm{CZ}_{\Delta_{i}}$ is the product of CZ gates along the three edges of the shaded face $\Delta_{i}$ to the right of the leftmost qubit $i$ and $\mathcal{V}=V\setminus\left(\mathcal{F}_{s}\cup\{i,f\}\right)$ is the set of all vertices in the bulk $(i,f)$ that are not part of a shaded face (i.e., the three middle dotted vertices in Fig. 6). Importantly, the endpoint operators for $\mathsf{S}_{x,t}$ and $\mathsf{S}_{z,t}$ both anticommute (on qubit $f$ and the shaded face $\Delta_{i}$), while the operators commute everywhere else (in the bulk).
The generators of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ are given by
$U(x)=\prod\limits_{j\notin\mathcal{F}_{s}}X_{j}~{},~{}~{}U(z)=\prod_{\Delta\in\mathcal{F}_{s}}X_{\Delta}\,,~{}~{}$ (75)
on an infinite chain, which are represented projectively by the endpoint operators in Eq. 6.2. The nontrivial string order of the hypergraph state is captured by
$\left\langle\Psi_{t}\middle|\,\mathsf{S}_{\nu,t}\middle|\Psi_{t}\right\rangle=1\,,~{}~{}$ (76)
for $\nu=x,z$. Note that any of the dotted vertices $j>i$ in Fig. 6 constitute valid right endpoints, and any tetrahedron corresponding to a dotted vertex $j<f$ and the shaded face $\Delta_{j}$ to its right constitutes a valid left endpoint. The chain in Fig. 6 can clearly be extended to be thermodynamically large, so that the hypergraph state has nontrivial SPT order protected by the symmetry group $\mathcal{G}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ generated by Eq. 75.
@@ 6.3 Cluster states for $k\geq 1$ logical qubits
Figure 7: Standard teleportation of $k=2$ logical qubits using a modification of the generalized cluster state $\lvert C_{2}\rangle$ (77). The conditioning of the recovery operations on the measurements is indicated by color-coded lines. The region within the dotted box is precisely the “bulk” resource state $\lvert C_{2}\rangle$ (77), where the operations outside the dotted box capture the required modifications near the boundaries.
We first consider an extension of the cluster-state teleportation protocol discussed in Secs. 2.4 and 6.1 Briegel and Raussendorf (2001) to $k\geq 1$ Stephen et al. (2022). We present a family of SPT states $\lvert C_{k}\rangle$ (77) with protecting symmetry $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ (as required by Theorem 18) that is closely related to other known families of resource states for $k$-qubit teleportation Raussendorf et al. (2022); Stephen et al. (2022).
@@
$\lvert C_{k}\rangle\equiv\bigg{(}\prod_{j}\mathrm{CZ}_{j,j+1}\,\prod_{j^{\prime}}H_{j^{\prime}}\bigg{)}^{k}\,\lvert y_{-}y_{-}\cdots y_{-}\rangle\,,~{}$ (77)
where $H_{j}=(X_{j}+Z_{j})/\sqrt{2}$ is the Hadamard gate, and the $+1$ eigenstates |+delimited-|$\lvert$+$\rangle$| + ⟩ of $X$ in Eq. 20 have been replaced by the $-1$ eigenstates $\lvert y_{-}\rangle$ of $Y$ (where $\lvert y_{\pm}\rangle=(\lvert 0\rangle\pm\mathrm{i}\,\lvert 1\rangle)/\sqrt{2}$ in the $Z$ basis). We note that the $k=1$ state $\lvert C_{1}\rangle$ (77) is unitarily equivalent to the standard cluster state Briegel and Raussendorf (2001) via a $Z$-basis rotation of all qubits by $\pi/4$. The stabilizer group $\operatorname*{stab}\left(*\right){\lvert C_{k}\rangle}$ is generated by the operators
$\mathsf{S}_{j}=Z_{j}\,Y_{j+1}\cdots Y_{j+2k-1}\,Z_{j+2k}\,,~{}~{}$ (78)
which bears some resemblance to Eq. 6.1. Furthermore, $\lvert C_{k}\rangle$ (77) has a $\mathbb{Z}_{2}^{2k}$ symmetry generated by
$g_{p}=\prod\limits_{j}\,Y_{2\,k\,j+p}\,,~{}~{}$ (79)
@@
Importantly, we note that the resource state $\lvert\Psi_{t}\rangle$ for teleportation requires modifying $\lvert C_{k}\rangle$ (77) in the vicinity of the initial and final logical sets $I$ and $F$. In the bulk of the chain (i.e., a distance $r>v\,T$ from both $I$ and $F$), the initial state $\lvert\Psi_{0}\rangle$ is simply $\lvert y_{-}\rangle$ on each site, and the unitary $\mathcal{U}$ is identical to the one implicit in Eq. 77. The required modifications are shown explicitly in Fig. 7 for $k=2$ logical qubits. They include modifying (i) the initial state $\lvert\Psi_{0}\rangle$ in $F$ according to $\lvert y_{-}\rangle\to\lvert+\rangle$; (ii) the unitary $\mathcal{U}$ in the region $I=\{1,2\}$ to include phase gates $S=\mathrm{diag}(1,\mathrm{i})$ in the computational basis; and (iii) the measurement channel $\mathcal{M}$ in the region $I$ to include $k=2$ additional measurements in $I$ (only the four measurements in the boxed “bulk” region in Fig. 7 repeat for $m>1$).
The important takeaway is that the resource state $\lvert\Psi_{t}\rangle$ (36) is equivalent to $\lvert C_{k}\rangle$ (77) in the bulk of the chain. By the same token, this state is compatible with Theorem 18 for any $k$, as the modifications to $\lvert C_{k}\rangle$ (77) required for teleportation are localized to the two edges, and merely modify the endpoint operators, which anticommute as required. Although it is possible to work out the modifications required at the two edges (i.e., outside of the boxed region in Fig. 7) for any $k>2$, this is not particularly illuminating or instructive. Instead, we consider an alternative family of states with a systematic procedure for encoding the edges in Sec. 6.4.

### v3

@@ 6.2 Hypergraph states
We present a $k=1$ standard teleportation protocol involving a hypergraph [52, 53, 25] resource state $\lvert\Psi^{\,}_{t}\rangle$ (36). Because $\operatorname*{stab}\left(*\right){\lvert\Psi^{\,}_{t}\rangle}\nsubseteq\mathbf{P}^{\vphantom{*}}_{*}{N}$ for the hypergraph state $\lvert\Psi^{\,}_{t}\rangle$, this example highlights how our results extend beyond “stabilizer states” [51]. The hypergraph state $\lvert\Psi^{\,}_{t}\rangle$ is defined on the vertices $v\in V$ of a graph $G=(V,E)$ comprising connected tetrahedra, as depicted in Fig. 6. The initial state is the same as for the cluster state (20)
$\lvert\Psi^{\vphantom{*}}_{0}\rangle=\lvert\psi\rangle^{\vphantom{\prime}}_{i}\otimes\lvert+++\cdots\rangle\otimes\lvert\boldsymbol{0}\rangle^{\vphantom{\prime}}_{\rm ss}\,,$ (68)
@@
$\mathcal{U}=\prod_{\Delta\in\mathcal{F}_{u}}\mathrm{CCZ}^{\vphantom{*}}_{\Delta}\,,$ (69)
where $\Delta$ runs over all unshaded faces (the set $\mathcal{F}_{u}$) of the graph in Fig. 6, and the controlled-controlled-$Z^{\vphantom{*}}$ gate $\mathrm{CCZ}^{\,}_{\Delta}$ applies CZ to two of the qubits on the face $\Delta$ if the third is in the $Z^{\vphantom{*}}$-basis state $\lvert 1\rangle$. In the computational ($Z^{\vphantom{*}}$) basis, $\mathrm{CCZ}^{\,}_{\Delta}=\mathrm{diag}(1,1,1,1,1,1,1,-1)$ sends $\lvert 111\rangle^{\,}_{\Delta}\to-\lvert 111\rangle^{\,}_{\Delta}$, and acts trivially otherwise.
We note that a length-three version of this hypergraph state appears in Ref. 53, but its connection to SPT order was not discussed. The hypergraph state in Fig. 6 can be viewed as a cluster state [43] whose odd sites correspond to the vertices connecting the tetrahedra (noted by dots in Fig. 6), and whose even sites are encoded in the three qubits around the shaded faces in Fig. 6.
The CCZ gate corresponds to the third order of the Clifford hierarchy [3], where the first order is the Pauli group $\mathbf{P}^{\vphantom{*}}_{*}{N}$, the second order is the Clifford group $\mathbf{C}^{\vphantom{*}}_{*}{N}$, which maps $\mathbf{P}^{\vphantom{*}}_{*}{N}$ to itself, while gates like CCZ map $\mathbf{C}^{\vphantom{*}}_{*}{N}$ to itself. In particular, for three vertices $i,j,k$, we have
$\mathrm{CCZ}^{\vphantom{*}}_{ijk}\,X^{\vphantom{*}}_{i}\,\mathrm{CCZ}^{\vphantom{*}}_{ijk}=X^{\vphantom{*}}_{i}\,\mathrm{CZ}^{\vphantom{*}}_{jk}\,,~{}$ (70)
@@
The hypergraph teleportation protocol $\mathcal{W}$ is straightforward. After applying the entangling unitary $\mathcal{U}$ (69) to the initial state $\lvert\Psi^{\,}_{0}\rangle$ (68), one next measures all qubits (except the rightmost site $f$) in the $X^{\vphantom{*}}$ basis in the channel $\mathcal{M}$. Thus, the measurement regions in Fig. 6 correspond to the tetrahedra formed by the shaded triangular faces and the vertex to the left. In the channel $\mathcal{R}$, the parity of the outcomes of all $X^{\vphantom{*}}_{j}$ measurements for qubits $j\in\mathcal{F}_{s}$ in the shaded triangles in Fig. 6 determine a final $\chi^{\,}_{z}=X^{\vphantom{*}}_{f}$ rotation, while the parity of the $X^{\vphantom{*}}_{j}$ outcomes for all other qubits $j\in V\setminus\mathcal{F}_{s}$ determine a final $\chi^{\,}_{x}=Z^{\vphantom{*}}_{f}$ rotation.
In the Heisenberg picture, the two logical operators $X^{\vphantom{*}}_{{\rm L}}$ and $Z^{\vphantom{*}}_{{\rm L}}$ evolve under the dilated channels $\mathcal{R}$ and $\mathcal{M}$ via
$X^{\prime}_{\rm L}$ $=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{M}^{\dagger}\,\mathcal{R}^{\dagger}\,X^{\vphantom{*}}_{f}\,\mathcal{R}\,\mathcal{M}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=\prod_{j\in V\setminus\mathcal{F}_{s}}X^{\vphantom{*}}_{j}$ (71a)
$Z^{\prime}_{\rm L}$ $=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{M}^{\dagger}\,\mathcal{R}^{\dagger}\,Z^{\vphantom{*}}_{f}\,\mathcal{R}\,\mathcal{M}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=Z^{\vphantom{*}}_{f}\prod_{\Delta\in\mathcal{F}_{s}}X^{\vphantom{*}}_{\Delta}\,,$ (71b)
acting on the hypergraph resource state $\lvert\Psi^{\,}_{t}\rangle$, where $\mathcal{F}_{s}$ denotes the shaded faces in Fig. 6, the set $V\setminus\mathcal{F}_{s}$ corresponds to the filled dots in Fig. 6, and $X^{\vphantom{*}}_{\Delta}=\prod_{j\in\Delta}\,X^{\vphantom{*}}_{j}$.
@@
$X^{\vphantom{*}}_{f}(T)$ $=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{W}^{\dagger}\,X^{\vphantom{*}}_{f}\,\mathcal{W}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=\prod_{j\in V\setminus\mathcal{F}_{s}}X^{\vphantom{*}}_{j}$ (72a)
$Z^{\vphantom{*}}_{f}(T)$ $=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{W}^{\dagger}\,Z^{\vphantom{*}}_{f}\,\mathcal{W}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=Z^{\vphantom{*}}_{i}\prod_{\Delta\in\mathcal{F}_{s}}X^{\vphantom{*}}_{\Delta}\,,$ (72b)
where $\mathcal{U}$ acts trivially on all $X^{\vphantom{*}}$ operators. Essentially, $\mathcal{U}$ (69) acts on $X^{\prime}_{\rm L}$ (71a) by attaching CZ operators to $X^{\vphantom{*}}_{j}$ along all edges of each shaded face whose vertices neighbor qubit $j$ (70). However, each shaded face borders two such $X^{\vphantom{*}}_{j}$ operators, and so each CZ operator is attached twice to $X^{\prime}_{\rm L}$ (71a), but $\mathrm{CZ}^{2}=\mathds{1}$. The action of $\mathcal{U}$ (69) on $Z^{\prime}_{\rm L}$ (71b) again follows from Eq. 70, and again, we find that CZ gates are always attached twice, so that the only change from Eq. 71b to Eq. 72b is that $Z^{\vphantom{*}}_{f}\to Z^{\vphantom{*}}_{i}$. Note that Eqs. 72 satisfy Prop. 1 for the initial state $\lvert\Psi^{\,}_{0}\rangle$ (68).
Finally, we identify the string order parameters establishing that the hypergraph resource state $\lvert\Psi^{\,}_{t}\rangle$ has nontrivial $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ SPT order. The string order parameters (53) correspond to the stabilizers $\mathsf{S}^{\,}_{\nu,t}$ given by
$\mathsf{S}^{\vphantom{\prime}}_{\nu,t}$ $=\mathcal{U}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}\,\mathcal{U}^{\dagger}=\mathcal{U}\,\sigma^{\nu}_{i}\,\sigma^{\nu}_{f}(T)\,\mathcal{U}^{\dagger}$
@@
which is the product of the initial-state logical operator $\sigma^{\nu}_{i}$ evolved “forward” under $\mathcal{U}$ (69) and the resource-state logical operators (71). We note that $Z^{\vphantom{*}}_{i}$ commutes with $\mathcal{U}$(69), and is thus unchanged, while $X^{\vphantom{*}}_{i}$ evolves according to Eq. 70. The resulting order parameters are
$\mathsf{S}^{\vphantom{\prime}}_{x,t}$ $=\mathrm{CZ}_{\Delta_{i}}\,\Big{(}\prod\limits_{j\in\mathcal{V}}\,X^{\vphantom{*}}_{j}\,\Big{)}\,X^{\vphantom{*}}_{f}$ (74a)
$\mathsf{S}^{\vphantom{\prime}}_{z,t}$ $=Z^{\vphantom{*}}_{i}\,X^{\vphantom{*}}_{\Delta_{i}}\,\Big{(}\prod\limits_{\begin{subarray}{c}\Delta\in\mathcal{F}_{s}\\ \delta\neq\Delta_{i}\end{subarray}}X^{\vphantom{*}}_{\Delta}\,\Big{)}\,Z^{\vphantom{*}}_{f}\,,~{}$ (74b)
where $\mathrm{CZ}_{\Delta_{i}}$ is the product of CZ gates along the three edges of the shaded face $\Delta_{i}$ to the right of the leftmost qubit $i$ and $\mathcal{V}=V\setminus\left(\mathcal{F}_{s}\cup\{i,f\}\right)$ is the set of all vertices in the bulk $(i,f)$ that are not part of a shaded face (i.e., the three middle dotted vertices in Fig. 6). Importantly, the endpoint operators for $\mathsf{S}^{\,}_{x,t}$ and $\mathsf{S}^{\,}_{z,t}$ both anticommute (on qubit $f$ and the shaded face $\Delta_{i}$), while the operators commute everywhere else (in the bulk).
The generators of the $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ are given by
$U(x)=\prod\limits_{j\notin\mathcal{F}_{s}}X^{\vphantom{*}}_{j}~{},~{}~{}U(z)=\prod_{\Delta\in\mathcal{F}_{s}}X^{\vphantom{*}}_{\Delta}\,,~{}~{}$ (75)
on an infinite chain, which are represented projectively by the endpoint operators in Eq. 74. The nontrivial string order of the hypergraph state is captured by
$\left\langle\Psi^{\,}_{t}\middle|\,\mathsf{S}^{\vphantom{\prime}}_{\nu,t}\middle|\Psi^{\,}_{t}\right\rangle=1\,,~{}~{}$ (76)
for $\nu=x,z$. Note that any of the dotted vertices $j>i$ in Fig. 6 constitute valid right endpoints, and any tetrahedron corresponding to a dotted vertex $j<f$ and the shaded face $\Delta_{j}$ to its right constitutes a valid left endpoint. The chain in Fig. 6 can clearly be extended to be thermodynamically large, so that the hypergraph state has nontrivial SPT order protected by the symmetry group $\mathcal{G}=\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ generated by Eq. 75.
@@ 6.3 Cluster states for $k\geq 1$ logical qubits
Figure 7: Standard teleportation of $k=2$ logical qubits using a modification of the generalized cluster state $\lvert C^{\,}_{2}\rangle$ (77). The conditioning of the recovery operations on the measurements is indicated by color-coded lines. The region within the dotted box is precisely the “bulk” resource state $\lvert C^{\,}_{2}\rangle$ (77), where the operations outside the dotted box capture the required modifications near the boundaries.
We first consider an extension of the cluster-state teleportation protocol discussed in Secs. 2.4 and 6.1 [43] to $k\geq 1$ [72]. We present a family of SPT states $\lvert C^{\,}_{k}\rangle$ (77) with protecting symmetry $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ (as required by Theorem 18) that is closely related to other known families of resource states for $k$-qubit teleportation [39, 72].
@@
$\lvert C^{\vphantom{\prime}}_{k}\rangle\equiv\bigg{(}\prod_{j}\mathrm{CZ}^{\,}_{j,j+1}\,\prod_{j^{\prime}}H^{\,}_{j^{\prime}}\bigg{)}^{k}\,\lvert y_{-}y_{-}\cdots y_{-}\rangle\,,~{}$ (77)
where $H^{\,}_{j}=(X^{\vphantom{*}}_{j}+Z^{\vphantom{*}}_{j})/\sqrt{2}$ is the Hadamard gate, and the $+1$ eigenstates |+delimited-|$\lvert$+$\rangle$| + ⟩ of $X^{\vphantom{*}}$ in Eq. 20 have been replaced by the $-1$ eigenstates $\lvert y_{-}\rangle$ of $Y^{\vphantom{*}}$ (where $\lvert y_{\pm}\rangle=(\lvert 0\rangle\pm\mathrm{i}\,\lvert 1\rangle)/\sqrt{2}$ in the $Z^{\vphantom{*}}$ basis). We note that the $k=1$ state $\lvert C^{\,}_{1}\rangle$ (77) is unitarily equivalent to the standard cluster state [43] via a $Z^{\vphantom{*}}$-basis rotation of all qubits by $\pi/4$. The stabilizer group $\operatorname*{stab}\left(*\right){\lvert C^{\,}_{k}\rangle}$ is generated by the operators
$\mathsf{S}^{\vphantom{\prime}}_{j}=Z^{\vphantom{*}}_{j}\,Y^{\vphantom{*}}_{j+1}\cdots Y^{\vphantom{*}}_{j+2k-1}\,Z^{\vphantom{*}}_{j+2k}\,,~{}~{}$ (78)
which bears some resemblance to Eq. 67. Furthermore, $\lvert C^{\,}_{k}\rangle$ (77) has a $\mathbb{Z}_{2}^{2k}$ symmetry generated by
$g^{\vphantom{\prime}}_{p}=\prod\limits_{j}\,Y^{\vphantom{*}}_{2\,k\,j+p}\,,~{}~{}$ (79)
@@
Importantly, we note that the resource state $\lvert\Psi^{\,}_{t}\rangle$ for teleportation requires modifying $\lvert C^{\,}_{k}\rangle$ (77) in the vicinity of the initial and final logical sets $I$ and $F$. In the bulk of the chain (i.e., a distance $r>v\,T$ from both $I$ and $F$), the initial state $\lvert\Psi^{\,}_{0}\rangle$ is simply $\lvert y_{-}\rangle$ on each site, and the unitary $\mathcal{U}$ is identical to the one implicit in Eq. 77. The required modifications are shown explicitly in Fig. 7 for $k=2$ logical qubits. They include modifying (i) the initial state $\lvert\Psi^{\,}_{0}\rangle$ in $F$ according to $\lvert y_{-}\rangle\to\lvert+\rangle$; (ii) the unitary $\mathcal{U}$ in the region $I=\{1,2\}$ to include phase gates $S=\mathrm{diag}(1,\mathrm{i})$ in the computational basis; and (iii) the measurement channel $\mathcal{M}$ in the region $I$ to include $k=2$ additional measurements in $I$ (only the four measurements in the boxed “bulk” region in Fig. 7 repeat for $m>1$).
The important takeaway is that the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) is equivalent to $\lvert C^{\,}_{k}\rangle$ (77) in the bulk of the chain. By the same token, this state is compatible with Theorem 18 for any $k$, as the modifications to $\lvert C^{\,}_{k}\rangle$ (77) required for teleportation are localized to the two edges, and merely modify the endpoint operators, which anticommute as required. Although it is possible to work out the modifications required at the two edges (i.e., outside of the boxed region in Fig. 7) for any $k>2$, this is not particularly illuminating or instructive. Instead, we consider an alternative family of states with a systematic procedure for encoding the edges in Sec. 6.4.